looked like a rectangle is 300 cm^2. the difference between length and width is 13 cm. find the lenght of the diagonal

Answer by checkley71(8403)

L=W+13
300=W(W+13)
300=W^2+13W
W^2+13W-300=0
(W-12)(W+25)
W-12=0
W=12cmNSWER FOR THE WIDTH.
L=12+13=25cm R THE LENGTH.
THE DIAGONAL IS C
A^2+B^2=C^2
12^2+25^2C^2
144+625=C^2
769=C^2
C=SQRT769
C=27.73cm IS THE LENGTH OF THE DIAGONAL.